OpenStudy (anonymous):

How many molecules are there in 15.7 liters of chlorine gas at STP?

OpenStudy (anonymous):

At STP 22.4 liter = 1 mole = 6.022 X 10 ^23 At STP 15.7 liter = 1/22.4 * 15.7 =0.701 * 6.022 X 10^23 = 4.22 X 10 ^23